# A Communication Consistent Approach to Signal Temporal Logic Task Decomposition in Multi-Agent Systems

### v2

@@ 5.1 Parametric tasks for decomposition
From the family of parametric predicate functions in (29) a decomposition of the inconsistent task $\varphi_{ij}$ is obtained by a two-step procedure: 1) define a path of agents $\bm{\pi}_{i}^{j}$ from $i$ to $j$ through the communication graph $\mathcal{G}_{c}$; 2) define a set of parametric tasks $\bar{\varphi}^{ij}_{rs}\;\forall(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ with corresponding predicate functions $\bar{h}^{ij}_{rs}$ as per (29) such that $(\bm{x}(t),0)\models\bar{\phi}^{ij}=\bigwedge_{(r,s)\in\epsilon(\bm{\pi}_{i}^{j})}\bar{\varphi}^{ij}_{rs}\Rightarrow(\bm{x}(t),0)\models\varphi_{ij}$. The solution to 1) is well studied in the literature so that we assume a path $\bm{\pi}_{i}^{j}$ between any pair $i,j\in\mathcal{V}$ can be found efficiently [39, Ch. 2]. On the other hand, a solution to step 2) is provided by Lemma 1.
@@
Thus, for every satisfying trajectory $(\bm{x}(t),0)\models\bar{\phi}^{ij}$ it must hold $\bm{e}_{ij}(\bar{t})\in\mathcal{B}_{ij}\Rightarrow h_{ij}(\bm{e}_{ij}(\bar{t}))\geq 0$. Since $\bar{t}\in[a,b]$, then by (15b) it holds $\rho^{\varphi_{ij}}(\bm{x}(t),0)=\min_{t\in[a,b]}(h_{ij}(\bm{e}_{ij}(t)))\geq 0\Rightarrow(\bm{x}(t),0)\models\varphi_{ij}$, which concludes the proof.
@@
Example 5.2.
Figure 4(a) shows a system of 6 drones with states $\bm{x}_{i}=\bm{p}_{i}$ being the positions of each drone, while on the left in Figure 4(b) the task graph $\mathcal{G}_{\psi}$ (dashed lines) and a communication graph $\mathcal{G}_{c}$ (solid lines) assigned to the agents are shown. The edge $(1,4)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ is an inconsistent edge that requires decomposition through $\bm{\pi}_{1}^{4}=[1,2,3,4]$ such that a new task $\bar{\phi}^{ij}=\bar{\varphi}_{12}^{14}\land\bar{\varphi}_{23}^{14}\land\bar{\varphi}_{34}^{14}$ is introduced as per Lemma 1. The new task graph resulting from the decomposition is then $\mathcal{G}_{\bar{\psi}}$ (on the right in Figure 4(b)). Figure 4(a) also shows the similar truth sets $\bar{\mathcal{B}}_{rs}^{14},\;\forall(r,s)\in\epsilon(\bm{\pi}_{1}^{4})=\{(1,2),(2,3),(3,4)\}$ (gray rectangles) and $\mathcal{B}_{14}$ (red rectangle). The inclusions relation (32) is graphically understood by letting the Minkowski sum $\bigoplus_{(r,s)\in\epsilon(\bm{\pi}_{1}^{4})}\bar{\mathcal{B}}_{rs}^{14}$ to be a subset of $\mathcal{B}_{14}$.
Figure 4: (a) A team of 6 drones with respective communication graph. (b)(Left) Initial task graph $\mathcal{G}_{\psi}$ (dashed line) and communication graph $\mathcal{G}_{c}$ (solid line) of the system. (Right) final task graph $\mathcal{G}_{\psi}$ after applying the proposed decomposition.
Note that when an inconsistent task $\varphi_{ij}$ featuring the eventually operator $F$ has to be decomposed, then the tasks $\bar{\varphi}_{rs}^{ij},\;\forall(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ are necessarily synchronized over a common time instant $\bar{t}$ as specified by (31). This synchronization is needed for the satisfaction of the original task $\varphi_{ij}$, although $\bar{t}$ can be chosen arbitrarily within $[a,b]$. Concerning the inclusions relation (32), it is known from Prop. 3 and Prop. 4 that (32) can be efficiently verified by a set of linear inequalities. Namely, consider again the truth set $\mathcal{B}_{ij}=\mathcal{P}(A_{ij},\bm{c}_{ij},\bm{z}_{ij})$ for the inconsistent task $\varphi_{ij}$ and let the decomposition path $\bm{\pi}_{i}^{j}$. Furthermore, let the similar polytopes $\bar{\mathcal{B}}_{rs}^{ij}=P(A_{ij},\bm{c}_{rs}^{ij},\alpha_{rs}^{ij}\bm{z}_{ij}),\;\forall(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ derived from the parametric task $\bar{\varphi}_{rs}^{ij}$ as per Lemma 1 with parameters $\bm{\eta}_{rs}^{ij}=[\bm{c}_{rs}^{ij},\alpha_{rs}^{ij}]$. By recalling the results of Prop. 3 and Prop. 4, and by letting the generator set $V_{ij}=\nu(\mathcal{P}(A_{ij},\bm{0},\bm{z}_{ij}))$, then the inclusion relation (32) is verified if it holds
@@
and $G_{k}=\left[I_{n}\mid\bm{v}_{k}\vphantom{X}\right]\in\mathbb{R}^{n\times(n+1)}$, $\bm{v}_{k}\in V_{ij},\,\forall k=1,\ldots|V_{ij}|$. From Prop. 5, if (32) holds, then also
@@
where (35) holds with equality if (32) does. We refer to the left-hand side of (35) as the decomposition accuracy of the decomposed task $\varphi_{ij}$. We show in Section 6 that maximising the decomposition accuracy on the right-hand side of (35) is the main objective of our decomposition approach
Hereafter, it is assumed that all the inconsistent tasks have been decomposed using a conjunction of parametric tasks as per Lemma 1. Next, in Section 5.2 and 5.3, we provide a set of constraints that are imposed over the parameters $\bm{\eta}_{rs}^{ij}$ such that the respective tasks $\bar{\varphi}_{rs}^{ij}$ are communication consistent and such that the tasks $\bar{\varphi}_{rs}^{ij}$ do not cause conflicts as per Facts 1-5.
5.2 Handling conflicting conjunctions
In this section, it is clarified how conflicting conjunctions as per Facts 1-5 might arise during decomposition and how to resolve them by appropriately constraining the parameters $\bm{\eta}_{rs}^{ij}$ associated with each $\bar{\varphi}_{rs}^{ij}$. Namely, we analyze how the consistent and conflict-free collaborative tasks $\phi_{rs}$ for all $(r,s)\in\mathcal{E}_{c}\cap\mathcal{E}_{\psi}$, that were in place before the decomposition, are modified by the introduction of the new tasks $\bar{\varphi}_{rs}^{ij},\;\forall(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ resulting from the decomposition, and how the consistency and conflict-free properties for these can be preserved. As in the previous sections, it is assumed that inconsistent tasks are without conjunctions as $\phi_{ij}=\varphi_{ij}$. A change of notation is required at this point of the presentation and applies hereafter. Namely, let the sets
@@ 5.2 Handling conflicting conjunctions
where the set $\mathcal{E}_{\pi}\subseteq\mathcal{E}_{c}$ contains the edges in $\mathcal{G}_{c}$ that are used for the decomposition of all the inconsistent tasks $\phi_{ij}$ with $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$, while $\Pi_{rs}\subseteq\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ contains the inconsistent edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ such that the decomposition path $\bm{\pi}_{i}^{j}$ is passing through $(r,s)\in\mathcal{E}_{\pi}\subseteq\mathcal{E}_{c}$. Moreover, let the collaborative task $\phi_{rs}:=\land_{k=1}^{K_{rs}}\varphi_{rs}^{k}$, as per (11c), be defined over $(r,s)\in\mathcal{E}_{\pi}$ before the decomposition. Then, after the decomposition of the inconsistent tasks $\varphi_{ij},\;\forall(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$, an additional conjunction of parametric tasks given by $\land_{(i,j)\in\Pi_{rs}}\bar{\varphi}_{rs}^{ij}$ is included over the edge $(r,s)\in\mathcal{E}_{\pi}$. Let then
@@
$\bigwedge_{(i,j)\in\Pi_{rs}}\bar{\varphi}^{ij}_{rs}\underset{y_{\Pi_{rs}}}{\equiv}\bigwedge_{k\in\bar{\mathcal{K}}_{rs}}\varphi^{k}_{rs},$ (39)
@@
where $\mathcal{K}_{rs}=\{1,\ldots K_{rs}\}\cup\bar{\mathcal{K}}_{rs}$. An example is provided to clarify the new notation.
Example 5.3.
@@
Consider Figure 5. The left panel represents the communication graph $\mathcal{G}_{c}$ (solid edges) together with the task graph $\mathcal{G}_{\psi}$ (dashed edges) such that tasks $\varphi_{24}$ and $\varphi_{25}$ are inconsistent. The consistent edge $(1,5)\in\mathcal{E}_{c}$ is on the decomposition paths $\bm{\pi}_{2}^{4}=[2,1,5,4]$ and $\bm{\pi}_{2}^{5}=[2,1,5]$ for both inconsistent tasks $\varphi_{24}$ and $\varphi_{25}$, while the task $\varphi_{15}$ is assigned before the decomposition. From the index mapping $y_{\Pi_{15}}$, with $y_{\Pi_{15}}((2,4))=2$ and $y_{\Pi_{15}}((2,5))=3$ the notational equivalence
$\varphi_{15}\land\bar{\varphi}_{15}^{24}\land\bar{\varphi}_{15}^{25}\underset{\text{$y_{\Pi_{rs}}(\cdot)$}}{\equiv}\varphi_{15}^{1}\land\varphi_{15}^{2}\land\varphi_{15}^{3},$
is established. The final task graph and communication graphs are then shown on the right panel in Figure 5.
@@
With this new notation, this section is concerned with stating sufficient conditions on the parameters $\bm{\eta}_{rs}^{k},\;\forall k\in\bar{\mathcal{K}}_{rs}$ associated with the parametric tasks $\varphi_{rs}^{k},\;\forall k\in\bar{\mathcal{K}}_{rs}$ such that the task $\phi_{rs}$ in (40) is not a conflicting conjunction. Indeed, we highlight again that after the introduction of the index map $y_{\Pi_{rs}}$, the tasks $\varphi_{rs}^{k},\;\forall k\in\bar{\mathcal{K}}_{rs}$ are parametric as per (31), with corresponding truth set given by $\mathcal{B}_{rs}^{k}=\mathcal{P}(A_{rs}^{k},\bm{c}_{rs}^{k},\alpha_{rs}^{k}\bm{z}^{k}_{rs})$ as per (29b) and such that $\bm{\eta}_{rs}^{k}:=[\bm{c}_{rs}^{k},\alpha_{rs}^{k}]$ is a parameter to be optimised. On the other hand, the tasks $\varphi_{rs}^{k},\;k\in\{1,\ldots K_{rs}\}$ are not parametric, as per (11b), and such that their truth set is a fixed polytope in the same general form $\mathcal{B}_{rs}^{k}=\mathcal{P}(A_{rs}^{k},\bm{c}_{rs}^{k},\alpha_{rs}^{k}\bm{z}^{k}_{rs})$, but with $\alpha_{rs}^{k}=1$ and $\bm{\eta}_{rs}^{k}=[\bm{c}_{rs}^{k},1]$ having fixed value as per (14). Note that, in general, the polytopes $\mathcal{B}_{rs}^{k},\forall k\in\mathcal{K}_{rs}$ are not similar (cf. Def 3). We are now ready to present the next main results. Following the presentation of the different types of conflicting conjunctions, we often refer to the set of indices $\mathcal{I}_{G}$ and $\mathcal{I}_{F}$ as per (23) where $\mathcal{I}_{G}\cup\mathcal{I}_{F}=\mathcal{K}_{rs}$, $\varphi^{k}_{rs}=G_{[a^{k},b^{k}]}\mu_{rs}^{k},\;\forall k\in\mathcal{I}_{G}$ and $\varphi^{k}_{rs}=F_{[a^{k},b^{k}]}\mu_{rs}^{k},\;\forall k\in\mathcal{I}_{F}$.
@@
Figure 5: Decomposition of formulas $\varphi_{25}$, $\varphi_{15}$, $\varphi_{24}$. (Left) Initial task graph (dashed) and communication graph (solid). (Right) Final task graph (dashed) and communication graph (solid).
Proposition 5.4.
@@
Consider $\phi_{rs}=\bigwedge_{k\in\mathcal{K}_{rs}}\varphi^{k}_{rs}$ as per (40) and the set of subset of indices $\mathfrak{L}_{rs}$ as per (24) over the edge $(r,s)$. Furthermore, let $\mathfrak{L}^{M}_{rs}\subseteq\mathfrak{L}_{rs}$ be defined as $\mathfrak{L}^{M}_{rs}=\{L\in\mathfrak{L}_{rs}\;|\;L\not\subset L^{\prime},\,\forall L^{\prime}\in\mathfrak{L}_{rs}\}$. Then if for each set of indices $L\in\mathfrak{L}^{M}_{rs}$ there exists $\bm{\xi}\in\mathbb{R}^{n}$, such that
@@
Proof 5.5.
By Prop. 4 we know that the set of inequalities (41) guarantees that for every $L\in\mathfrak{L}^{M}_{rs}$ it holds $\bigcap_{l\in L}\mathcal{B}_{rs}^{l}\supseteq\{\bm{\xi}\}\neq\emptyset$, thus disproving the conditions for conflicting conjunction over the sets of indices $L_{rs}\in\mathfrak{L}^{M}_{rs}$ as per Fact 1. Consider now a set $L^{\prime}$ such that $L^{\prime}\in\mathfrak{L}_{rs}\setminus\mathfrak{L}^{M}_{rs}$. Then necessarily, by definition of $\mathfrak{L}^{M}_{rs}$, there must be an $L\in\mathfrak{L}^{M}_{rs}$ such that $L^{\prime}\subset L$. We can then write $\bigcap_{l\in L^{\prime}}\mathcal{B}_{rs}^{l}\supseteq\bigcap_{l\in L}\mathcal{B}_{rs}^{l}\supseteq\{\bm{\xi}\}\neq\emptyset$. Since $L^{\prime}$ was chosen arbitrarily from $\mathfrak{L}_{rs}$, then the conditions for conflicting conjunction as per Facts 1 do not hold for any $L^{\prime}\in\mathfrak{L}_{rs}$ concluding the proof.
The set $\mathfrak{L}^{M}_{rs}$ represents the set of subsets in $\mathfrak{L}_{rs}$ that are “maximal”, meaning that any set of indices $L\in\mathfrak{L}^{M}_{rs}$ is not contained in any other set $L^{\prime}\in\mathfrak{L}_{rs}$. Therefore the result of Prop. 5.4 suggests that checking for conflicting conjunctions over the subset of indices in $\mathfrak{L}^{M}_{rs}$ is sufficient to avoid conflicting conjunctions of type 1, instead of checking all subsets in $\mathfrak{L}_{rs}$ where $|\mathfrak{L}^{M}_{rs}|\leq|\mathfrak{L}_{rs}|$. Second, once the power set of indices $2^{\mathcal{I}_{G}}$ is computed, then extracting the set $\mathfrak{L}_{rs}\subseteq 2^{\mathcal{I}_{G}}$ (and thus $\mathfrak{L}^{M}_{rs}$) is computationally efficient as it requires only to verify the condition $\max_{l\in L}a^{l}\leq\min_{l\in L}b^{l}$ for each set $L\in 2^{\mathcal{I}_{G}}$, with $a^{l},b^{l}]$ being the time interval of the respective tasks $\varphi_{rs}^{l}$. Third, only the sets of indices $L\in\mathfrak{L}^{M}_{rs}$ such that $L\cap\bar{\mathcal{K}}_{rs}\neq\emptyset$ have to be checked for conflicting conjunctions since, by Assumption 2, the conjunction $\land_{l\in\{1,\ldots K_{rs}\}}\varphi_{rs}^{l}$ as per (40) is not a conflict of type 1.
Proposition 5.6.

### v3

@@ V-A Parametric tasks for decomposition
From the family of parametric predicate functions in (30) a decomposition of the inconsistent task $\varphi_{ij}$ is obtained by a two-step procedure: 1) define a path of agents $\bm{\pi}_{i}^{j}$ from $i$ to $j$ through the communication graph $\mathcal{G}_{c}$; 2) define a set of parametric tasks $\bar{\varphi}^{ij}_{rs}\;\forall(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ with corresponding predicate functions $\bar{h}^{ij}_{rs}$ as per (30) such that $(\bm{x}(t),0)\models\bar{\phi}^{ij}=\bigwedge_{(r,s)\in\epsilon(\bm{\pi}_{i}^{j})}\bar{\varphi}^{ij}_{rs}\Rightarrow(\bm{x}(t),0)\models\varphi_{ij}$. The solution to 1) is well studied in the literature so that we assume a path $\bm{\pi}_{i}^{j}$ between any pair $i,j\in\mathcal{V}$ can be found efficiently [40, Ch. 2]. On the other hand, a solution to step 2) is provided by Lemma 1.
@@
Thus, for every satisfying trajectory $(\bm{x}(t),0)\models\bar{\phi}^{ij}$ it must hold $\bm{e}_{ij}(\bar{t})\in\mathcal{B}_{ij}\Rightarrow h_{ij}(\bm{e}_{ij}(\bar{t}))>0$. Since $\bar{t}\in[a,b]$, then by (16b) it holds $\rho^{\varphi^{ij}}(\bm{x}(t),0)=\min_{t\in[a,b]}(h_{ij}(\bm{e}_{ij}(t)))>0\Rightarrow(\bm{x}(t),0)\models\varphi_{ij}$, which concludes the proof. ∎
@@
Example 3
Figure 4(a) shows a system of 6 drones with states $\bm{x}_{i}=\bm{p}_{i}$ being the positions of each drone, while Figure 4(b) shows the task graph $\mathcal{G}_{\psi}$ and a communication graph $\mathcal{G}_{c}$ assigned to the agents. The edge $(1,4)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ is an inconsistent edge that requires decomposition through $\bm{\pi}_{1}^{4}=[1,2,3,4]$ such that a new task $\bar{\phi}^{ij}=\bar{\varphi}_{12}^{14}\land\bar{\varphi}_{23}^{14}\land\bar{\varphi}_{34}^{14}$ is introduced as per Lemma 1. The new task graph resulting from the decomposition is then $\mathcal{G}_{\bar{\psi}}$ (bottom in Figure 4(b)). Figure 4(a) also shows the similar truth sets $\bar{\mathcal{B}}_{rs}^{14},\;\forall(r,s)\in\epsilon(\bm{\pi}_{1}^{4})=\{(1,2),(2,3),(3,4)\}$ (gray rectangles) and $\mathcal{B}_{14}$ (red rectangle). The inclusions relation (33) is graphically understood by letting the Minkowski sum $\bigoplus_{(r,s)\in\epsilon(\bm{\pi}_{1}^{4})}\bar{\mathcal{B}}_{rs}^{14}$ to be a subset of $\mathcal{B}_{14}$.
Figure 4: Graphical representation of a task decomposition process according to Sec. V-A for a team of 6 drones. Detailed explanation is provided in Example 3.
Note that when an inconsistent task $\varphi_{ij}$ featuring the eventually operator $F$ has to be decomposed, then the tasks $\bar{\varphi}_{rs}^{ij},\;\forall(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ are necessarily synchronized over a common time instant $\bar{t}$ as specified by (32). This synchronization is needed for the satisfaction of the original task $\varphi_{ij}$, although $\bar{t}$ can be chosen arbitrarily within $[a,b]$. Concerning the inclusions relation (33), it is known from Prop. 3 and Prop. 4 that (33) can be efficiently verified by a set of linear inequalities. Namely, consider again the truth set $\mathcal{B}_{ij}=\mathcal{P}(A_{ij},\bm{c}_{ij},\bm{z}_{ij})$ for the inconsistent task $\varphi_{ij}$ and let the decomposition path $\bm{\pi}_{i}^{j}$. Furthermore, let the similar polytopes $\bar{\mathcal{B}}_{rs}^{ij}=P(A_{ij},\bm{c}_{rs}^{ij},\alpha_{rs}^{ij}\bm{z}_{ij}),\;\forall(r,s)\in\epsilon(\bm{\pi}_{i}^{j})$ derived from the parametric task $\bar{\varphi}_{rs}^{ij}$ as per Lemma 1 with parameters $\bm{\eta}_{rs}^{ij}=[\bm{c}_{rs}^{ij},\alpha_{rs}^{ij}]$. If we let the generator set $V_{ij}=\nu(\mathcal{P}(A_{ij},\bm{0},\bm{z}_{ij}))$, then the inclusion relation (33) is verified if it holds
@@
and $G_{k}=\left[I_{n}\mid\bm{v}_{k}\vphantom{X}\right]\in\mathbb{R}^{n\times(n+1)}$, $\bm{v}_{k}\in V_{ij},\,\forall k=1,\ldots|V_{ij}|$. At the same time, it is known from Prop. 5 that if (33) holds, then the satisfaction of (33) implies
@@
where (36) holds with equality if (33) does. We refer to the left-hand side of (36) as the decomposition accuracy of the decomposed task $\varphi_{ij}$. We show in Section VI that maximising the decomposition accuracy on the right-hand side of (36) is indeed the main objective of our decomposition approach. Next, in Section V-C and V-B, we provide a set of constraints that are imposed over the parameters $\bm{\eta}_{rs}^{ij}$ such that the respective tasks $\bar{\varphi}_{rs}^{ij}$ are communication consistent and such that the tasks $\bar{\varphi}_{rs}^{ij}$ do not cause conflicts as per Facts 1-5.
V-B Handling conflicting conjunctions during decomposition
In this section, it is clarified how conflicting conjunctions as per Facts 1-5 might arise during decomposition and how to resolve them by appropriately constraining the parameters $\bm{\eta}_{rs}^{ij}$ associated with each $\bar{\varphi}_{rs}^{ij}$. Namely, we analyze how the consistent and conflict-free collaborative tasks $\phi_{rs}$ for all $(r,s)\in\mathcal{E}_{\pi}\subseteq\mathcal{E}_{c}$, that were in place before the decomposition, are modified by the introduction of the new tasks $\bar{\varphi}_{rs}^{ij},\;\forall(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ resulting from the decomposition, and how the consistency and conflict-free properties for these can be preserved. As in the previous sections, it is assumed that inconsistent tasks are without conjunctions as $\phi_{ij}=\varphi_{ij}$. A change of notation is required at this point of the presentation and applies hereafter. Namely, let the sets
@@ V-B Handling conflicting conjunctions during decomposition
where the set $\mathcal{E}_{\pi}\subseteq\mathcal{E}_{c}$ contains all the edges in $\mathcal{G}_{c}$ that are used for the decomposition of all the inconsistent tasks $\phi_{ij}$ with $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$, while $\Pi_{rs}\subseteq\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ contains all the inconsistent edges $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ such that the decomposition path $\bm{\pi}_{i}^{j}$ is passing though $(r,s)\in\mathcal{E}_{\pi}\subseteq\mathcal{E}_{c}$. Moreover, let the collaborative task $\phi_{rs}:=\land_{k=1}^{K_{rs}}\varphi_{rs}^{k}$, as per (12c), be defined over $(r,s)\in\mathcal{E}_{\pi}$ before the decomposition. Then, after the decomposition of all the inconsistent tasks $\varphi_{ij},\;\forall(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$, we have that an additional conjunction of parametric tasks given by $\land_{(i,j)\in\Pi_{rs}}\bar{\varphi}_{rs}^{ij}$ is included over the edge $(r,s)\in\mathcal{E}_{\pi}$. Let then
@@
$\bigwedge_{(i,j)\in\Pi_{rs}}\bar{\varphi}^{ij}_{rs}=\bigwedge_{k\in\bar{\mathcal{K}}_{rs}}\varphi^{k}_{rs},$ (40)
@@
where $\mathcal{K}_{rs}=\{1,\ldots K_{rs}\}\cup\bar{\mathcal{K}}_{rs}$. A clarifying example is provided.
Example 4
@@
Consider Figure 5. The upper panel represents the communication graph (solid edges) together with the task graph (dashed edges) such that tasks $\varphi_{58},\;\varphi_{57}$ and $\varphi_{51}$ are inconsistent. The consistent edge $(6,7)\in\mathcal{E}_{c}$ is on the decomposition paths $\bm{\pi}_{5}^{8}=[5,6,7,8]$ and $\bm{\pi}_{5}^{7}=[5,6,7]$ for both inconsistent tasks $\varphi_{58}$ and $\varphi_{57}$, while the task $\phi_{67}=\varphi_{67}$ is assigned to the edge before the decomposition. From the index mapping $y_{\Pi_{67}}$, with $y_{\Pi_{67}}((5,7))=2$ and $y_{\Pi_{67}}((5,8))=3$ it derives
$\varphi_{67}\land\varphi_{67}^{57}\land\varphi_{67}^{58}\underset{\text{$y_{\Pi_{rs}}(\cdot)$}}{=}\varphi_{67}^{1}\land\varphi_{67}^{2}\land\varphi_{67}^{3},$
where the index $1$ was assigned to the first task for consistency with the fact that $K_{rs}=1$. The final task graph and communication graphs are then represented by the lower panel in Figure 5.
@@
With this new notation, this section is concerned with stating sufficient conditions on the parameters $\bm{\eta}_{rs}^{k},\;\forall k\in\bar{\mathcal{K}}_{rs}$ associated with the parametric tasks $\varphi_{rs}^{k},\;\forall k\in\bar{\mathcal{K}}_{rs}$ such that the task $\phi_{rs}$ in (41) is not a conflicting conjunction. Indeed, we highlight again that after the introduction of the index map $y_{\Pi_{rs}}$, the tasks $\varphi_{rs}^{k},\;\forall k\in\bar{\mathcal{K}}_{rs}$ are as per (32), with corresponding truth set given by $\mathcal{B}_{rs}^{k}=\mathcal{P}(A_{rs}^{k},\bm{c}_{rs}^{k},\alpha_{rs}^{k}\bm{z}^{k}_{rs})$ as per (30b) and such that $\bm{\eta}_{rs}^{k}:=[\bm{c}_{rs}^{k},\alpha_{rs}^{k}]$ is a parameter to be optimised. On the other hand, the tasks $\varphi_{rs}^{k},\;k\in\{1,\ldots K_{rs}\}$ are not parametric, as per (12b), and such that their truth set is a fixed polytope in the same general form $\mathcal{B}_{rs}^{k}=\mathcal{P}(A_{rs}^{k},\bm{c}_{rs}^{k},\alpha_{rs}^{k}\bm{z}^{k}_{rs})$, but with $\alpha_{rs}^{k}=1$ and $\bm{\eta}_{rs}^{k}=[\bm{c}_{rs}^{k},1]$ having fixed value as per (15). Note that, in general, the polytopes $\mathcal{B}_{rs}^{k},\forall k\in\mathcal{K}_{rs}$ are not similar (cf. Def 3). We are now ready to present the next main results. Following the presentation of the different types of conflicting conjunctions, we often refer to the set of indices $\mathcal{I}_{G}$ and $\mathcal{I}_{F}$ as per (24) where $\mathcal{I}_{G}\cup\mathcal{I}_{F}=\mathcal{K}_{rs}$, $\varphi^{k}_{rs}=G_{[a^{k},b^{k}]}\mu_{rs}^{k},\;\forall k\in\mathcal{I}_{G}$ and $\varphi^{k}_{rs}=F_{[a^{k},b^{k}]}\mu_{rs}^{k},\;\forall k\in\mathcal{I}_{F}$.
@@
Figure 5: Graphical decomposition of formulas $\varphi_{58}$, $\varphi_{57}$, $\varphi_{51}$. (Top) Initial task graph (dashed lines) and communication graph (solid). (Bottom) Final task graph (dashed lines) and communication graph (solid).
Proposition 6
@@
Consider $\phi_{rs}=\bigwedge_{k\in\mathcal{K}_{rs}}\varphi^{k}_{rs}$ as per (41) and the set of subset of indices $\mathfrak{L}_{rs}$ as per (25) over the edge $(r,s)$. Furthermore, let $\mathfrak{L}^{M}_{rs}\subseteq\mathfrak{L}_{rs}$ be defined as $\mathfrak{L}^{M}_{rs}=\{L\in\mathfrak{L}_{rs}\;|\;L\not\subset L^{\prime},\,\forall L^{\prime}\in\mathfrak{L}\}$. Then if for each set of indices $L\in\mathfrak{L}^{M}_{rs}$ there exists $\bm{\xi}\in\mathbb{R}^{n}$, such that
@@
Proof:
By Prop. 4 we know that the set of inequalities (42) guarantees that for every $L\in\mathfrak{L}^{M}_{rs}$ it holds $\bigcap_{l\in L}\mathcal{B}_{rs}^{l}\supseteq\{\bm{\xi}\}\neq\emptyset$, thus disproving the conditions for conflicting conjunction over the sets of indices $L_{rs}\in\mathfrak{L}^{M}_{rs}$ as per Fact 1. Consider now a set $L^{\prime}$ such that $L^{\prime}\in\mathfrak{L}_{rs}\setminus\mathfrak{L}^{M}_{rs}$. Then necessarily, by definition of $\mathfrak{L}^{M}_{rs}$, there must be an $L\in\mathfrak{L}^{M}_{rs}$ such that $L^{\prime}\subset L$. We can then write $\bigcap_{l\in L^{\prime}}\mathcal{B}_{rs}^{l}\supseteq\bigcap_{l\in L}\mathcal{B}_{rs}^{l}\supseteq\{\bm{\xi}\}\neq\emptyset$. Since $L^{\prime}$ was chosen arbitrarily from $\mathfrak{L}_{rs}$, then the conditions for conflicting conjunction as per Facts 1 do not hold for any $L^{\prime}\in\mathfrak{L}_{rs}$ concluding the proof. ∎
The set $\mathfrak{L}^{M}_{rs}$ represents the set of subsets in $\mathfrak{L}_{rs}$ that are “maximal”, in the sense that any set of indices $L\in\mathfrak{L}^{M}_{rs}$ is not contained in any other set $L^{\prime}\in\mathfrak{L}_{rs}$. Therefore the result of Prop. 6 suggests that checking for conflicting conjunctions over the subset of indices in $\mathfrak{L}^{M}_{rs}$ is sufficient to avoid conflicting conjunctions of type 1, instead of checking all the possible subset of indices in $\mathfrak{L}_{rs}$ where $|\mathfrak{L}^{M}_{rs}|\leq|\mathfrak{L}_{rs}|$. Second, once the power set of indices $2^{\mathcal{I}_{G}}$ is computed, then extracting the set $\mathfrak{L}_{rs}\subseteq 2^{\mathcal{I}_{G}}$ (and thus $\mathfrak{L}^{M}_{rs}$) is computationally efficient as it requires only to verify the condition $\max_{l\in L}a^{l}\leq\min_{l\in L}b^{l}$ for each set $L\in 2^{\mathcal{I}_{G}}$. Third, only the sets of indices $L\in\mathfrak{L}^{M}_{rs}$ such that $L\cap\bar{\mathcal{K}}_{rs}\neq\emptyset$ have to be checked for conflicting conjunctions since, by Assumption 2, the conjunction $\land_{l\in\{1,\ldots K_{rs}\}}\varphi_{rs}^{l}$ as per (41) is not a conflict of type 1.
Proposition 7